OpenStudy (anonymous):

remove the parenthesis and simplify 3a+5b-8(5a-7b)

OpenStudy (anonymous):

\[3a+5b-8(5a-7b)\] To get rid of the parentheses, we need to use distributive property. So, \[-8\times5a=-40a\] \[-8\times -7b=56b\] Which turns the original expression into; \[3a+5b-40a+56b\] And now all that's left is to combine like terms \[3a-40a=-37a\] \[5b+56b=61b\] So that forms; \[-37a+61b\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem.